# Egocentric Action-aware Inertial Localization in Point Clouds

### v1

@@ Abstract
This paper presents a novel inertial localization framework named Egocentric Action-aware Inertial Localization (EAIL), which leverages egocentric action cues from head-mounted IMU signals to localize the target individual within a 3D point cloud. Human inertial localization is challenging due to IMU sensor noise that causes trajectory drift over time. The diversity of human actions further complicates IMU signal processing by introducing various motion patterns. Nevertheless, we observe that some actions observed through the head-mounted IMU correlate with spatial environmental structures (e.g., bending down to look inside an oven, washing dishes next to a sink), thereby serving as spatial anchors to compensate for the localization drift. The proposed EAIL framework learns such correlations via hierarchical multi-modal alignment. By assuming that the 3D point cloud of the environment is available, it contrastively learns modality encoders that align short-term egocentric action cues in IMU signals with local environmental features in the point cloud. These encoders are then used in reasoning the IMU data and the point cloud over time and space to perform inertial localization. Interestingly, these encoders can further be utilized to recognize the corresponding sequence of actions as a by-product. Extensive experiments demonstrate the effectiveness of the proposed framework over state-of-the-art inertial localization and inertial action recognition baselines. Project page: https://github.com/mf-zhang/Ego-Inertial-Localization.
@@ 1 Introduction
Inertial localization has emerged as a pivotal technology for human tracking across a wide range of applications, including personal navigation [22] and augmented reality [52]. Inertial measurement units (IMUs) are commonly embedded in wearable devices such as smartphones, smart glasses [12], and headsets [46]. When affixed to human bodies, IMUs can capture acceleration and angular velocity to record 3D human movements. Compared to vision-based localization methods [40, 28], inertial localization enables user tracking in an energy-efficient and privacy-preserving manner [30].
@@
The key technical contribution is a hierarchical modality alignment technique that effectively extracts the correlation between the action cues and the point cloud. Specifically, we first learn the encoders for short-term IMU signals and local point clouds to be aligned with the corresponding egocentric point-of-view images and textual descriptions of observed actions in a contrastive fashion. The learned IMU and point-cloud encoders are then invoked in spatio-temporal reasoning modules to jointly predict possible actions occurring at each moment and the locations in the global point cloud where the predicted actions can be observed. While our primary goal is the localization in the latter part, the predicted locations can also improve inertial action recognition as a beneficial by-product.
@@ 5.2 Inertial Localization Results
{tblr}
colspec=X[1,l] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c], cell11 = r=2l, cell13 = c=4c, cell18 = c=4c, hline1,9 = -0.08em, hline2 = 3-6, hline2 = 8-11, hline3 = -, width=1.1 Method & Seen Rooms Unseen Rooms
@@
MnasNet[58] 2.95 7.57 12.75 / 2.16 5.45 9.86 /
EfficientNet[57] 3.1 7.77 12.71 / 2.54 6.68 10.84 /
IMUNet[67] 3.74 10.15 17.23 / 3.4 9.17 14.63 /
RoNIN[22] 4.86 12.77 20.65 / 3.96 9.52 15.65 /
NILoc[24]$+$ 17.03 41.31 74.15 88.17 13.32 37.85 69.21 84.08
Ours 43.86 70.15 89.6 96.01 26.86 65.97 90.79 89.55
@@ 5.3 Inertial Action Recognition Results
{tblr}
colspec=X[1,l] X[0.1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c], cell11 = r=2c, cell13 = c=2c, cell16 = c=2c, hline1,7 = -0.08em, hline2 = 3-4, hline2 = 6-7, hline3 = -, width=1.25 Method & Seen Rooms Unseen Rooms
@@
DeepConvLSTM[45] 15.20 43.27 12.47 36.86
EVIMAE[70]-- 18.34 46.12 9.30 31.10
IMU2CLIP[42] 18.96 50.43 12.27 37.04
@@
{tblr}
colspec=X[2.25,l] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c], cell11 = r=2l, cell13 = c=6c, cell110 = c=6c, cell31 = c=15l,fg=myblue!50, cell61 = c=15l,fg=myblue!50, cell91 = c=15l,fg=myblue!50, hline1 = -0.08em, hline2 = 3-8, hline2 = 10-15, hline3 = -0.08em, hline6 = -0.08em, hline9 = -0.08em, hline13 = -0.08em, width=1.2 Method & Seen Rooms Unseen Rooms
@@
w/o vision language 39.75 66.30 87.56 95.70 13.21 51.02 20.41 61.95 89.83 88.13 9.53 29.09
w/ vision language 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
@@
w/o action loss 41.92 68.54 87.75 95.51 / / 25.37 63.79 89.28 89.02 / /
w/ action loss 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
@@ 5.4 Ablation Studies
{tblr}
colspec=X[1,l] X[0.1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c], cell11 = r=2c, cell13 = c=2c, cell16 = c=2c, hline1,6 = -0.08em, hline2 = 3-4, hline2 = 6-7, hline3 = -, width=1.25 Method & Seen Rooms Unseen Rooms
@@ 7 Conclusion
We introduce a novel framework EAIL for inertial localization that uniquely leverages egocentric action cues from head-mounted IMU signals to enhance localization within 3D point clouds. By aligning short-term IMU signals with local point cloud structures, and further integrating temporal and spatial reasoning, our approach effectively addresses common challenges like trajectory drift and complex human actions. Additionally, our method demonstrates a robust capacity for both trajectory prediction and action recognition across diverse settings, as validated by rigorous evaluations on the EgoExo4D dataset. This work underscores the potential for action-environment correlations to serve as powerful anchors in practical inertial localization tasks, opening avenues for future explorations into more advanced sensing and positioning technologies.
@@
{tblr}
colspec=X[2.25,l] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c], cell11 = r=2l, cell13 = c=6c, cell110 = c=6c, cell31 = c=15l,fg=myblue!50, cell71 = c=15l,fg=myblue!50, cell111 = c=15l,fg=myblue!50, cell151 = c=15l,fg=myblue!50, hline1 = -0.08em, hline2 = 3-8, hline2 = 10-15, hline3 = -0.08em, hline7 = -0.08em, hline11 = -0.08em, hline15 = -0.08em, hline18 = -0.08em, width=1.1 Method & Seen Rooms Unseen Rooms

### v2

@@ Abstract
This paper presents a novel inertial localization framework named Egocentric Action-aware Inertial Localization (EAIL), which leverages egocentric action cues from head-mounted IMU signals to localize the target individual within a 3D point cloud. Human inertial localization is challenging due to IMU sensor noise that causes trajectory drift over time. The diversity of human actions further complicates IMU signal processing by introducing various motion patterns. Nevertheless, we observe that some actions captured by the head-mounted IMU correlate with spatial environmental structures (e.g., bending down to look inside an oven, washing dishes next to a sink), thereby serving as spatial anchors to compensate for the localization drift. The proposed EAIL framework learns such correlations via hierarchical multi-modal alignment with vision-language guidance. By assuming that the 3D point cloud of the environment is available, it contrastively learns modality encoders that align short-term egocentric action cues in IMU signals with local environmental features in the point cloud. The learning process is enhanced using concurrently collected vision and language signals to improve multimodal alignment. The learned encoders are then used in reasoning the IMU data and the point cloud over time and space to perform inertial localization. Interestingly, these encoders can further be utilized to recognize the corresponding sequence of actions as a by-product. Extensive experiments demonstrate the effectiveness of the proposed framework over state-of-the-art inertial localization and inertial action recognition baselines. Project page: https://github.com/mf-zhang/Ego-Inertial-Localization.
@@ 1 Introduction
Inertial localization has emerged as a pivotal technology for human tracking across a wide range of various applications, including personal navigation [22] and augmented reality [51]. Inertial measurement units (IMUs) are commonly embedded in wearable devices such as smartphones, smart glasses [12], and headsets [45]. When affixed to human bodies, IMUs can capture acceleration and angular velocity to record 3D human movements. Compared to vision-based localization methods [39, 28], inertial localization enables user tracking in an energy-efficient and privacy-preserving manner.
@@
The key technical contribution is a hierarchical modality alignment technique that effectively extracts the correlation between the action cues and the point cloud with vision-language guidance. Specifically, we first learn the encoders for short-term IMU signals and local point clouds to be aligned with the corresponding egocentric point-of-view images and textual descriptions of observed actions in a contrastive fashion. By incorporating guidance from pretrained vision-language models, we achieve improved multimodal feature alignment. The learned IMU and point-cloud encoders are then invoked in spatio-temporal reasoning modules to jointly predict possible actions occurring at each moment and the locations in the global point cloud where the predicted actions can be observed. While our primary goal is the localization in the latter part, the predicted locations can also improve inertial action recognition as a beneficial by-product.
@@ 5.2 Inertial Localization Results
Method $SeenRooms$ Unseen Rooms
@@
MnasNet[57] 2.952.952.95 7.57 12.75 / 2.16 5.45 9.86 /
EfficientNet[56] 3.103.103.10 7.77 12.71 / 2.54 6.68 10.84 /
IMUNet[66] 3.743.743.74 10.15 17.23 / 3.40 9.17 14.63 /
RoNIN[22] 4.864.864.86 12.77 20.65 / 3.96 9.52 15.65 /
NILoc[24]$+$ 17.0317.0317.03 41.31 74.15 88.17 13.32 37.85 69.21 84.08
Ours 43.86 70.15 89.60 96.01 26.86 65.97 90.79 89.55
@@ 5.3 Inertial Action Recognition Results
Method $SeenRooms$ Unseen Rooms
@@
DeepConvLSTM[44] 15.2015.2015.20 43.27 12.47 36.86
EVIMAE[69]-- 18.3418.3418.34 46.12 9.30 31.10
IMU2CLIP[41] 18.9618.9618.96 50.43 12.27 37.04
@@
Method $SeenRooms$ Unseen Rooms
@@
w/o vision language 39.7539.7539.75 66.30 87.56 95.70 13.21 51.02 20.41 61.95 89.83 88.13 9.53 29.09
w/ vision language 43.86 70.15 89.60 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
@@
w/o action loss 41.9241.9241.92 68.54 87.75 95.51 / / 25.37 63.79 89.28 89.02 / /
w/ action loss 43.86 70.15 89.60 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
@@ 5.4 Ablation Studies
Method $SeenRooms$ Unseen Rooms
@@ 7 Conclusion
We propose EAIL, a novel framework for inertial localization that uses egocentric action cues from head-mounted IMU signals to improve positioning in 3D point clouds. By aligning short-term IMU data with local structures and incorporating temporal-spatial reasoning, EAIL addresses challenges like trajectory drift and complex actions. Evaluated on the EgoExo4D dataset, it shows strong performance in trajectory prediction and action recognition. This highlights the potential of action-environment correlations as anchors for inertial localization, paving the way for advanced sensing and positioning methods.
8 Acknowledgement
This work was supported by JSPS KAKENHI Grant Number JP24K02956, JST ASPIRE Grant Number JPMJAP2303, JST SPRING Grant Number JPMJSP2108.
@@ 8 Acknowledgement
Method $SeenRooms$ Unseen Rooms